# Spatioformer: A Geo-encoded Transformer for Large-Scale Plant Species Richness Prediction

### v2

@@ I Introduction
In this study, we aim to predict the spatiotemporal distribution of plant species richness in Australia from EO imagery with geolocation context being taken into account. Considering that the relationship between plant species richness and remote sensing imagery varies from one location to another due to differences in vegetation composition, we propose Spatioformer, where a novel geolocation encoder is coupled with the transformer model in order to incorporate the geolocation context. The performance of Spatioformer in richness mapping is compared with a CNN model, a ViT model, and the FactoFormer model where the geolocational information is not encoded. Through quantitative analyses, we seek to address primarily the following important questions: (1) Does Spatioformer perform better than state-of-the-art algorithms in predicting plant species richness over large spatial scales? (2) What are the spatial patterns of plant species richness in Australia inferred from remote sensing evidence? (3) Where future in-situ surveys should be conducted as suggested by the mapping results?
@@ III-B Geolocation Encoder
A graphic illustration of the proposed geolocation encoder is shown in Fig. 3 for two example locations. The geolocation encoding vectors for these two locations are constructed with values referenced from the corresponding positions on a total of $d$ encoding layers. This geolocation encoder could provide geo-gradient information not only in the cardinal directions but also in various diagonal directions (as shown in Fig. 6 in Subsection III-B). The $d$, $a$, and $c$ values configured in this study for plant species richness mapping are described in Subsection IV-B.
@@ III-C The Spatioformer Structure
where $\lambda$ is a learnable parameter balancing the relative contribution between pixel values and geolocational information, given that these two different types of data may differ in their magnitudes. The geolocation-encoded embeddings $\bm{e}^{\prime}_{i}$ are then fed into the transformer encoder, together with a learnable geolocation-independent token at the pixel/patch level to account for geolocation-independent components in the input-output relationship. A fully connected layer is set as the output layer to produce the predicted value or class, depending on whether the problem to be solved is a regression or classification problem. The configuration of Spatioformer for plant species richness mapping is described in Subsection IV-B. The geolocation encoder in Spatioformer enables incorporation of geolocation context into the images. Therefore, this model , with the geolocation context being leveraged in the modelling.
@@ III-E Plant Species Richness Mapping
In this study, the Spatioformer model was trained with ground-truth richness samples and their corresponding Landsat images. With the trained model, plant species richness maps were compiled through inferring richness values from Landsat images. Mapping results were demonstrated for the years from 2015 to 2023. Compared with older years, maps for these recent years can better inform current and future conservation activities and policy making. We then aggregated these annual maps to compile the mean richness map across the six years in order to study plant species richness with enhanced spatial patterns. We also calculated the standard deviation map of richness over the six years to identify places where predicted plant species richness showed high temporal variations. The size of input images was ranged from 1 × 1 to 9 × 9 pixels in order to examine how mapping accuracy varied when different spatial scales of plant community were taken into account.
@@ IV-B Spatioformer Model Setup and Hyperparameter Selection
The input images of 9 × 9 pixels, with each pixel being of 30 m × 30 m spatial resolution, were embedded into a 16-dimensional feature space using a linear feed-forward layer. A geolocation token of 16 dimensions was added to the embedding of each pixel. The geolocation-encoded pixel embeddings were fed into three layers of eight-head self-attention module with each layer followed by a 64-dimensional feed-forward network. A fully connected network with a hidden layer of 1024 nodes was set as the output layer. The mean squared error between model-predicted and ground-truth richness values was employed as the loss function. The Adam optimizer was employed for network optimisation.
@@ IV-C Benchmarking with State-of-the-Art Models
The performance of Spatioformer was compared with a CNN model [58], a ViT model [25], and the FactoFormer model [30]. The CNN model consisted of three convolutional layers, with each layer followed by a ReLU activation layer and a batch normalisation layer. Each convolutional layer consisted of eight filters with a kernel size of 3 × 3. Following the convolutional layers were a flatten layer and a fully connected network. For the ViT model, each pixel was first embedded into a 16-dimensional feature space via a linear feed-forward layer, which was then fed into three layers of eight-head self-attention module. Each self-attention layer was followed by a feed-forward network of 64 dimensions. A flatten layer and a fully connected network were set as the output layers. For the FactoFormer model, each input image cube was first split into nonoverlapping tokenized patches along spectral and spatial dimensions, followed by processing them with two transformers simultaneously. Then the outputs of each transformer were flatten, concatenated, and passed to a a fully connected network.
@@ V-C Model Performance
Fig. 8 shows the predicted distribution of plant species richness across Australia in the year 2020, generated by the three benchmark models: CNN [58] (Fig. 8a), ViT [25] (Fig. 8b), and FactoFormer [30] (Fig. 8c). The relative difference of these maps to the map produced by the proposed Spatioformer model (Fig. 7f) reveals notable discrepancies in certain regions, as shown in Fig. 8d–f. For instance, the Jarrah Forest bioregion near the southwestern coast of Western Australia is a biodiverse region with richness generally higher than 60 species per 400 m2, as evidenced by the ground-truth samples (Fig. 2a). While the high richness in this region was successfully mapped with the proposed Spatioformer model (Fig. 7f), the three benchmark models predicted this region to have 30 $\sim$ 40 species per 400 m2 (Fig. 8a–c), roughly 20 $\sim$ 40 lower relative to the Spatioformer predictions (Fig. 8d–f). Another example is the Nullarbor Plain along the northwest coast of the Great Australian Bight. The extremely dry climate in this region has resulted in poor plant diversity of less than 10 species per 400 m2, as supported by field surveys shown in Fig. 2a. While the Spatioformer model accurately predicted the very low richness values observed in the Nullarbor Plain (Fig. 7f), the predicted richness values by three benchmark models (Fig. 8a–c) were not as low as those indicated by the ground-truth samples, showing about 10 $\sim$ 20 higher relative to the Spatioformer map (Fig. 8d–f). These results suggested that, the proposed Spatioformer model is capable of effectively adapting to local richness characteristics, thanks to the incorporation of geolocation context in its modelling.
As regions with a similar level of richness may present quite different spectral features in remote sensing imagery, predicting plant species richness over large spatial scales is a challenging task. Previous studies suggested that the relationship between plant species richness and remote sensing observations is location-specific, primarily due to difference in the assemblages of plant species among regions that results from their unique biogeographic histories, climate conditions, and geographical features [20]. For example, statistical analysis showed that the Southern Tablelands and Snowy Mountains regions differ in their relationships between spectral data and plant species richness [13]. Therefore, geolocation context might be useful for mapping richness across regions. The location-dependent modelling with Spatioformer encoded such geolocational information into the input spectral data, which could help accommodate the location dependency in mapping richness over large spatial scales. The results shown in Table I, and the visual comparison between Fig. 8 and Fig. 7f, indicated that it is beneficial to utilise geolocation context in large-scale richness mapping, with a higher overall accuracy being observed.
@@
Figure 8: Predicted plant species richness distribution in Australia with (a) Convolutional Neural Network (CNN) [58], (b) Vision Transformer (ViT) [25], and (c) FactoFormer [30], for the year 2020, and (d–f) their relative difference to the result obtained with the proposed Spatioformer for the same year (Fig. 7f).
@@ V-E Uncertainty in Richness Prediction
It is worth noting that several factors could contribute to the uncertainty in richness predictions. Areas with sparse ground sampling, such as the interior of Australia, exhibited higher prediction uncertainties, highlighting the model’s sensitivity to the availability of ground-truth data. Moreover, the quality of in-situ samples used for model training may affect the accuracy of the predictions. The model’s performance is another critical factor, as its ability to generalize across diverse landscapes would also influence the reliability of the predicted richness values.

### v3

@@ I Introduction
In this study, we aim to predict the spatiotemporal distribution of plant species richness in Australia from EO imagery with geolocation context being taken into account. Considering that the relationship between plant species richness and remote sensing imagery varies from one location to another due to differences in vegetation composition, we propose Spatioformer, where a novel geolocation encoder is coupled with the transformer model in order to incorporate the geolocation context. The performance of Spatioformer in richness mapping is compared with a CNN model, a ViT model, and the FactoFormer model where the geolocational information is not encoded. Through quantitative analyses, we seek to address primarily the following important questions: (1) Does Spatioformer perform better than state-of-the-art algorithms in predicting plant species richness over large spatial scales? (2) What are the spatial patterns of plant species richness in Australia inferred from remote sensing evidence? (3) Where should future in-situ surveys be conducted as suggested by the mapping results?
@@ III-B Geolocation Encoder
A graphic illustration of the proposed geolocation encoder is shown in Fig. 3 for two example locations. The geolocation encoding vectors for these two locations are constructed with values referenced from the corresponding positions in a total of $d$ encoding layers. This geolocation encoder could provide geo-gradient information not only in the cardinal directions but also in various diagonal directions (as shown in Fig. 6 in Subsection III-B). The $d$, $a$, and $c$ values configured in this study for plant species richness mapping are described in Subsection IV-B.
@@ III-C The Spatioformer Structure
where $\lambda$ is a learnable parameter balancing the relative contribution between pixel values and geolocational information, given that these two different types of data may differ in their magnitudes. The geolocation-encoded embeddings $\bm{e}^{\prime}_{i}$ are then fed into the transformer encoder, together with a learnable geolocation-independent token at the pixel/patch level to account for geolocation-independent components in the input-output relationship. A fully connected layer is set as the output layer to produce the predicted value or class, depending on whether the problem to be solved is a regression or classification problem. The configuration of Spatioformer for plant species richness mapping is described in Subsection IV-B. The geolocation encoder in Spatioformer integrates geolocation context into the images, enabling the model to leverage this information in modelling plant species richness distribution in this study.
@@ III-E Plant Species Richness Mapping
In this study, the Spatioformer model was trained with ground-truth richness samples and their corresponding Landsat images. With the trained model, plant species richness maps were compiled through inferring richness values from Landsat images. Mapping results were demonstrated for the years from 2015 to 2023. Compared with older years, maps for these recent years can better inform current and future conservation activities and policy making. We then aggregated these annual maps to compile the mean richness map across the nine years in order to study plant species richness with enhanced spatial patterns. We also calculated the standard deviation map of richness over the nine years to identify places where predicted plant species richness showed high temporal variations. The size of input images was ranged from 1 × 1 to 9 × 9 pixels in order to examine how mapping accuracy varied when different spatial scales of plant community were taken into account.
@@ IV-B Spatioformer Model Setup and Hyperparameter Selection
The input images of 9 × 9 pixels, with each pixel being of 30 m × 30 m spatial resolution, were embedded into a 16-dimensional feature space using a linear feed-forward layer. In cases where input images have irregular edges, a patching strategy can be implemented by applying zero-padding along the edges to standardise the input dimensions. A geolocation token of 16 dimensions was added to the embedding of each pixel, where the embedding feature dimension of 16 was determined based on multiple test runs. The geolocation-encoded pixel embeddings were fed into three layers of eight-head self-attention module with each layer followed by a 64-dimensional feed-forward network. A fully connected network with a hidden layer of 1024 nodes was set as the output layer. The mean squared error between model-predicted and ground-truth richness values was employed as the loss function. The Adam optimizer was employed for network optimisation.
@@ IV-C Benchmarking with State-of-the-Art Models
The performance of Spatioformer was compared with a CNN model [58], a ViT model [25], and the FactoFormer model [30]. The CNN model consisted of three convolutional layers, with each layer followed by a ReLU activation layer and a batch normalisation layer. Each convolutional layer consisted of eight filters with a kernel size of 3 × 3. Following the convolutional layers were a flatten layer and a fully connected network. For the ViT model, each pixel was first embedded into a 16-dimensional feature space via a linear feed-forward layer, which was then fed into three layers of eight-head self-attention module. Each self-attention layer was followed by a feed-forward network of 64 dimensions. A flatten layer and a fully connected network were set as the output layers. For the FactoFormer model, each input image cube was first split into non-overlapping tokenized patches along spectral and spatial dimensions, followed by processing them with two transformers simultaneously. Then the outputs of each transformer were flatten, concatenated, and passed to a a fully connected network.
@@ V-C Model Performance
Fig. 8 shows the predicted distribution of plant species richness across Australia in the year 2020, generated by the three benchmark models: CNN [58] (Fig. 8a), ViT [25] (Fig. 8b), and FactoFormer [30] (Fig. 8c). The relative difference of these maps to the map produced by the proposed Spatioformer model (Fig. 7f) reveals notable discrepancies in certain regions, as shown in Fig. 8d–f. For instance, the Jarrah Forest bioregion near the southwestern coast of Western Australia is a biodiverse region with richness generally higher than 60 species per 400 m2, as evidenced by the ground-truth samples (Fig. 2a). While the high richness in this region was successfully mapped with the proposed Spatioformer model (Fig. 7f), the three benchmark models predicted this region to have 30 $\sim$ 40 species per 400 m2 (Fig. 8a–c), roughly 20 $\sim$ 40 lower relative to the Spatioformer predictions (Fig. 8d–f). The difference in richness prediction results between Spatioformer and the benchmark models is shown in more detail in Fig. 9 with enlarged maps of the predicted plant species richness distribution in the Jarrah Forest bioregion. As observed from Fig. 9, the proposed Spatioformer model predicted the richness to be generally higher than 60 species per 400 m2 (Fig. 9d), matching better with the ground-truth evidence (Fig. 2a) than the benchmark models which estimated roughly 30 to 40 species per 400 m2 (Fig. 9a–c). Another example is the Nullarbor Plain along the northwest coast of the Great Australian Bight. The extremely dry climate in this region has resulted in poor plant diversity of less than 10 species per 400 m2, as supported by field surveys shown in Fig. 2a. While the Spatioformer model accurately predicted the very low richness values observed in the Nullarbor Plain (Fig. 7f), the predicted richness values by three benchmark models (Fig. 8a–c) were not as low as those indicated by the ground-truth samples, showing about 10 $\sim$ 20 higher relative to the Spatioformer map (Fig. 8d–f). These results suggested that, the proposed Spatioformer model is capable of effectively adapting to local richness characteristics, thanks to the incorporation of geolocation context in its modelling.
@@
Figure 8: Predicted plant species richness distribution in Australia with (a) Convolutional Neural Network (CNN) [58], (b) Vision Transformer (ViT) [25], and (c) FactoFormer [30], for the year 2020, and (d–f) their relative difference to the result obtained with the proposed Spatioformer for the same year (Fig. 7f). Richness values are in unit of number of species per 400 m2.
Figure 9: Predicted plant species richness distribution in the Jarrah Forest bioregion near the southwestern coast of Western Australia with (a) Convolutional Neural Network (CNN) [58], (b) Vision Transformer (ViT) [25], (c) FactoFormer [30], and (d) the proposed Spatioformer for the year 2020. Richness values are in unit of number of species per 400 m2.
As regions with a similar level of richness may present quite different spectral features in remote sensing imagery, predicting plant species richness over large spatial scales is a challenging task. Previous studies suggested that the relationship between plant species richness and remote sensing observations is location-specific, primarily due to difference in the assemblages of plant species among regions that results from their unique biogeographic histories, climate conditions, and geographical features [20]. For example, statistical analysis showed that the Southern Tablelands and Snowy Mountains regions differ in their relationships between spectral data and plant species richness [13]. Therefore, geolocation context might be useful for mapping richness across regions. The location-dependent modelling with Spatioformer encoded such geolocational information into the input spectral data, which could help accommodate the location dependency in mapping richness over large spatial scales. The results shown in Table I, the visual comparison between Fig. 8 and Fig. 7f, and the enlarged comparison maps shown in Fig. 9, indicated that it is beneficial to utilise geolocation context in large-scale richness mapping, with a higher overall accuracy being observed.
@@ V-E Uncertainty in Richness Prediction
It is worth noting that several factors could contribute to the uncertainty in richness predictions. Areas with sparse ground sampling, such as the interior of Australia, exhibited higher prediction uncertainties, highlighting the model’s sensitivity to the availability of ground-truth data. Moreover, the quality of in-situ samples used for model training may affect the accuracy of the predictions. The model’s performance is another critical factor, as its ability to generalise across diverse landscapes would also influence the reliability of the predicted richness values.